OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x - y = 10

OpenStudy (anonymous):

solve for y just in the first equation ? Not in the system of equations ?

OpenStudy (anonymous):

yeah thats what it is asking

OpenStudy (anonymous):

x + 3y = 6 3y = -x + 6 y = (-x + 6)/3 or y = -1/3x + 2

OpenStudy (anonymous):

Solve 7x - 2y = -3 14x + y = 14 so here i add the like terms?

OpenStudy (anonymous):

7x+13x?

OpenStudy (anonymous):

what are you trying to find ?

OpenStudy (anonymous):

Is it the same as the first?

OpenStudy (anonymous):

it just says solve

OpenStudy (anonymous):

Solve 7x - 2y = -3 14x + y = 14

OpenStudy (anonymous):

substitution or elimination ?

OpenStudy (anonymous):

which one is better?

OpenStudy (anonymous):

substitution is easier...we will do that. 14x + y = 14 (we need to isolate a variable) y = -14x + 14 now we sub -14x + 14 in for y in the other equation 7x - 2y = -3 7x - 2(-14x + 14) = -3 (distribute through the parenthesis) 7x + 28x - 28 = -3 (add 28 to both sides) 7x + 28x = -3 + 28 (combine like terms) 35x = 25 x = 25/35 x = 5/7 now we sub 5/7 in for x y = -14x + 14 y = -14(5/7) + 14 y = - 70/7 + 14 y = - 10 + 14 y = 4 check.. 7x - 2y = -3 7(5/7) - 2(4) = -3 35/7 - 8 = -3 5 - 8 = -3 -3 = -3 (correct) x = 5/7 and y = 4

OpenStudy (anonymous):

do you have any questions at all ?

OpenStudy (anonymous):

no i got every step, thank you..

OpenStudy (anonymous):

your very welcome :)

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

sure...what you got ?